# thanks: Present Address: Imec, Kapeldreef 75, 3001 Leuven, Belgium

### v1

@@ Abstract
SUPPLEMENTAL NOTE 3: Magnetic hysteresis of the resistance pattern
@@
Supplemental Figure S4: Resistance dependent on magnetic field for: (a) the main text 4TJJ array, (b) the second 4TJJ array, (c) and the reference two-terminal Josephson junction. The blue lines indicate a magnetic sweep direction from negative to positive and the orange lines from positive to negative. All devices show hysteretic behavior with magnetic field. The two upper resistance oscillations stem from the frustration pattern of the array, the lower resistance oscillation stems from the superconducting quantum interference of the two-terminal Josephson junction.
SUPPLEMENTAL NOTE 4: Determination of the resistance oscillation period
For the determination of the oscillation frequencies, fast Fourier transforms have been applied to the data in Fig. 3 c) and d) of the main text (see Fig. S5). The frequency spectra of both data curves match. The dominant frequency (period) is 0.16 1/mT (6.25 mT). The period of the beating pattern cannot be seen in the spectrum, probably because it is too close to the zero peak.
Supplemental Figure S5: Fast Fourier transforms of the data curves in main text Fig. 3 c) (fine sweep) and d) (wide range). The dominant frequency (period) is 0.16 1/mT (6.25 mT). The other peaks of the spectrum are higher harmonics of that frequency.
SUPPLEMENTAL NOTE 5: Description of the array device
This section gives a detailed description of the array device in the main text, including geometrical properties of the array lattice (see Fig. S6). According to figure S6 a), the unit cell of the array is a square of size $A_{\text{UC}}=573\;\text{nm}\cdot 567\;\text{nm}=(570\;\text{nm})^{2}$. The geometry of the weak link area between two neighboring leads of one 4TJJ (see Fig. S6 b) is described, on average, by the following dimensions: The width of the weak link is 129 nm and its length is 97 nm. The 4TJJ leads have a length of 471 nm.
During fabrication of the device, a trench is etched into the SiO2 (300 nm) / Si2N3 (100 nm) layer, down to the Si(111) substrate. On the substrate, the device is fabricated and, afterwards, the whole chip is covered with superconducting material. To minimize disturbance effects, the superconductor around the device is etched away as a last fabrication step, leaving an edge of superconductor of approximately 1 µm around the trench behind. To prevent this closed loop of superconductor from generating screening currents, the edge is disconnected close to the ends of the device contacts (see figure S6 c-d).
@@
SUPPLEMENTAL NOTE 6: Characteristic RCSJ values between two leads of a 4TJJ in the arrray
In this section, relevant parameters for the theoretical description of the array are derived from geometrical considerations of the prior section, as well as from measurement data of the reference two-terminal Josephson junction. We assume that critical current density and resistivity, derived in section 2, are geometry independent. We then use these values, together with the 4TJJ geometry to derive capacitance, resistivity, and critical current between two leads of one 4TJJ.
The coplanar capacitance is calculated using the geometrical values given above. The relative permeability $\epsilon_{r}$ is assumed to be one, since the weak link is a metal. This results in a capacitance of $C=2.729\cdot 10^{-18}$ F. The critical current between two 4TJJ leads is $I_{\mathrm{C}}=7.9$ µA. This results in a Josephson energy of $E_{\mathrm{J}}=\Phi_{0}I_{\mathrm{c}}/2\pi=16.2$ meV. The resistance is equal to $R=9.1\;\Omega$.
SUPPLEMENTAL NOTE 7: Additional measurements of the main text 30$\times$30 4TJJ array
@@
Supplemental Figure S7: IV measurements at 80 mK. a) The DC IV curve with a critical current of $I_{\mathrm{c}}=57\;\mu$A. b) AC differential resistance measured with the lock-in technique. The AC bias was 200 nA.
@@
Figure S9 shows the same resistance oscillations with magnetic field as before but, here, they were measured with the AC lock-in technique and no DC bias. The resulting resistance is lower than with the DC measurement because the time-dependent sinusoidal signal includes the superconducting state of the device as well. When comparing figure S9 a) with figure S4 a), one can see that the oscillations are identical in magnetic field, although the peak pattern is more symmetrical with the AC method.
@@
Figure LABEL:fig:SUPP_MISC_RB_DC_I_T a) shows resistance measurements dependent on magnetic field and DC current. At higher currents, superconductivity is not realized at any magnetic field and the oscillation amplitudes decline.
Figure LABEL:fig:SUPP_MISC_RB_DC_I_T b) shows resistance measurements dependent on magnetic field and temperature. At higher temperatures, superconductivity is not realized at any magnetic field and the oscillation amplitudes decline.
@@
SUPPLEMENTAL NOTE 8: Details of the theoretical modelling
@@
We here summarize the model calculations for the theoretical discussion of the main text. We describe the multiterminal Josephson junction array in the overdamped regime (discarding capacitive contributions), we get at each node $j$ the equation of motion,
$\frac{1}{2eR}\sum_{\langle j^{\prime}\rangle_{j}}\left(\dot{\phi}_{j}-\dot{\phi}_{j^{\prime}}\right)=-I_{c}\sum_{\langle j^{\prime}\rangle_{j}}\sin(\phi_{j}-\phi_{j^{\prime}}+a_{jj^{\prime}})+\sum_{\langle j^{\prime}\rangle_{j}}\eta_{jj^{\prime}}(t)\ ,$ (S1)
with shunt resistance $R$, critical current $I_{c}$, and the current noise term $\eta_{jj^{\prime}}$ due to thermal fluctuations from the resistor connecting nodes $j$ and $j^{\prime}$, where
$\langle\eta_{ii^{\prime}}(t)\eta_{jj^{\prime}}(t^{\prime})\rangle=\frac{2k_{B}T}{R}\delta_{ij}\delta_{i^{\prime}j^{\prime}}\delta(t-t^{\prime})\ .$ (S2)

### v2

@@ Abstract
Figure 2: Measurement data of the $30\times 30$ 4TJJ array at 80 mK. a) Differential resistance as a function of bias current and magnetic field. A periodic oscillation of the critical current is clearly visible. b) Resistance as a function of magnetic field with an applied dc bias of 30 $\mu$A. The resistance oscillations correspond to a periodicity of 6.25 mT. At around ±$\pm 5\;f_{\mathrm{uc}}$, the resistance oscillations are damped. c) Resistance oscillations under magnetic field ranging to ±$\pm 140$ mT with an applied dc bias of 30 $\mu$A. In total, 30 flux quantum oscillations are present.
Theoretical discussion. To explain the measured behavior of the 4TJJ array, we introduce a corresponding theoretical model. Based on the resistively capacitively shunted junction (RCSJ) network model for multi-terminal Josephson junctions [29, 30, 37, 39], the array can be described, as shown in Fig. 1 a), with four 2TJJs creating one 4TJJ [55]. This introduces a second lattice of areas pierced by magnetic flux, the central weak link region of the 4TJJs. The multi-terminal array model can therefore be conveniently represented in terms of an ordinary $J_{x}\times J_{y}$ square lattice junction model [where $J_{x}$ and $J_{y}$ are the horizontal and vertical number of superconducting nodes, see Fig. 1 b)], with the following important difference to previous theoretical and experimental studies: instead of all the plaquettes being pierced by the same flux, we get an alternating (checker board) flux pattern ($f=\Phi/\Phi_{0}$ and $f^{\prime}=\Phi^{\prime}/\Phi_{0}$). We deploy a classical RCSJ model approach, where the equations of motion for the superconducting phases at all nodes follow as usual from the Kirchhoff laws [55]. We focus on the overdamped regime  [^1] Using the extracted values for individual junction capacitances, resistances, and critical currents, we estimate the Steward-McCumber parameter to be $\approx 5.4\times 10^{-6}$, well within the overdamped regime., neglecting the capacitive contributions to the equations of motion, and assume zero temperature. As will become apparent below, it is instructive to include a finite Fraunhofer effect [14] for the individual junctions in the lattice.
Including a source and drain contact on two sides of the lattice (across which the bias current $I$ is applied), the RCSJ model allows for a direct computation of the dc resistance $R_{\text{dc}}$ as a function of the magnetic field. We do so by an explicit numerical evaluation of the classical RCSJ equations of motion  [^2] We use the NDSolve routine on Mathematica for explicit evaluation of the equations of mtion.. The results of the simulation are summarized in Fig. 3. For computational simplicity, the calculations were performed on a small array of $J_{x}=J_{y}=6$.
It is instructive to first consider the special case $f,f^{\prime}\in\mathbb{Z}$ (equivalent to $f=f^{\prime}=0$). Here, the dc resistance can be found analytically (as both quantum and thermal phase slips are absent) [55],
$\frac{R_{\text{dc}}}{R}=(J_{x}+1)\sqrt{\frac{1}{J_{y}^{2}}-\frac{I_{c}^{2}}{I^{2}}}\ ,$ (1)
where $R$ and $I_{c}$ are the individual junction resistance and critical current, respectively. The array thus transitions from superconducting ($R_{\text{dc}}=0$) to resistive ($R_{\text{dc}}>0$) when the bias current exceeds $J_{y}I_{c}$. For a regular square lattice [$f=f^{\prime}$, see also Fig. 3 b)] it is well known that the array generally leaves the superconducting regime for noninteger values of $f$, even though $I<J_{y}I_{c}$. For sufficiently low temperatures and current biases, the dc resistance nonetheless experiences dips at special rational values – known as the frustration pattern.
Consider now the general checker board array, $f\neq f^{\prime}$ [left panel in Fig. 3 b)], where we denote the (constant) ratio $\beta=f/f^{\prime}>1$. Here, there emerges a beating pattern, with the two characteristic periods $\beta>1$ and $\beta/(1+\beta)<1$ [see Figs. 3 a) and c)]. The larger period ($\beta$) is in general only approximate; e.g., for incommensurate $\beta$ the system can only approximately return to mutually integer $f,f^{\prime}$ – exhibiting a quasiperiodicity in flux space. The irrational nature of $\beta$ is however also highly relevant for the smaller period (which is exact): it separates the points where the sum of two neighbouring plaquette fluxes is zero, i.e., $f+f^{\prime}\in\mathbb{Z}$ [see top inset in Fig. 3 a)]. For the experimentally extracted value of $\beta\approx 10.9$ (cf. Fig. 2[55]) the difference between the two frequencies is large, leading to a clearly visible beating pattern [Fig. 3 c)]. Once $\beta$ is known, the central weak link area of a 4TJJ can be determined via $A^{\prime}=A_{\text{uc}}/(1+\beta)$, leading to $A^{\prime}\approx(165\;\text{nm})^{2}$ [55]. Crucially, for integer $f+f^{\prime}$, the repartition of the total flux into the two neighbouring plaquettes ($f$ and $f^{\prime}$) does not need to occur for special (integer or rational) value, since the ratio of the two fluxes, $\beta$, is in general incommesurate. Consequently, the system exhibits a stabilization of superconducting (BKT) phase (dips in $R_{\text{dc}}$) even for irrational $f,f^{\prime}$, a feature we choose to name ’frustrated frustration’.
Figure 3: Results of the theoretical analysis. In all curves (a,c-g), the $x$-axis is $f$, and the $y$-axis is $R_{\text{dc}}/R$. (a) $R_{\text{dc}}$ as a function of $f$ for the checker board model with $f/f^{\prime}=10.9$ and $I/I_{c}=5.0$. (b) Checker board ($f\neq f^{\prime}$) versus regular ($f=f^{\prime}$) lattice models. (c) Zoomed out version of (a) showing the beating pattern. (d) $R_{\text{dc}}$ for the regular lattice model $f=f^{\prime}$, with all other parameters the same as in (a,c). (e,f) $R_{\text{dc}}$ including Fraunhofer effect (red curves, see main text), where for the checker board model $f_{0}=30$ (e) and for the regular lattice $f_{0}=10.9$ (f). (g) $R_{\text{dc}}$ for $f/f^{\prime}=10.9$ (no Fraunhofer) for decreasing bias current. Data shifted for clarification. (h) Equilibrium loop current configuration ($I=0$) for different values of $f,f^{\prime}$. The dots indicate counter- (red) or clockwise (blue) going currents and their size represents the magnitude of the loop current (relative linear scale).
We now include the Fraunhofer effect, $I_{c}\sim\text{sinc}(\pi f/f_{0})$, where the parameter $f_{0}$ captures the junction area. The resulting reduction of the critical current at finite $f$ leads to a base offset in the $R_{\text{dc}}$-curve, see Fig. 3 e) [the red curve represents Eq. (1) with $f$-dependent $I_{c}$], in remarkable resemblance to the experimental data, Fig. 2 c). As a comparison, take the regular square lattice, $f=f^{\prime}$, where a qualitatively very similar beating pattern arises due to Fraunhofer [Fig. 3 d) versus Fig. 3 f)], by setting $f_{0}$ (instead of $f/f^{\prime}$) to $\approx 10.9$. However, this alternative model can be safely excluded to explain the experimentally observed phenomenology. First, the Fraunhofer beating pattern in Fig. 3 f) skips a beat at $f=0$, due to the sinc-function having no zero at the origin, such that there are two beating pattern frequencies, $2/f_{0}$ (main beat around $f=0$) and $1/f_{0}$ (side beats) [see Fig. 3 f)]. Figure 2 only exhibits one beating frequency, in alignment with the checker board model 3 c). Moreover, the area ratio extracted from micrographs of the device align much better with the first model ($f/f^{\prime}\approx 10.9$ and much larger $\sim f_{0}$). Related to that, for $f/f_{0}\ll 1$ (while $I\gg JI_{c}$), the base offset is approximately linear, $R_{\text{dc}}\sim f$, consistent with the measured asymptotics in Fig. 2. Indeed, nodes separating two beats [green dots in Figs. 3 e) and f)] can only exhibit an increase in $R_{\text{dc}}$ in the checker board model, while in the regular lattice, the onset of the beating pattern must coincide with $R_{\text{dc}}$ reaching the constant plateau value $\approx R$.
We observe that the pattern in Fig. 3 a) transitions from resistance dips to peaks at integer $f+f^{\prime}$, depending on whether $f$ and $f^{\prime}$ are individually closer to integer or half-integer. Those peaks can become dips when lowering the bias current, see Fig. 3 g), marking the onset of the checker board version of regular frustration – which goes hand in hand with the formation of additional resistance dips at $f+f^{\prime}$ half integer [see, $10<f<$ 1 in Fig. 3 g)]  [^3] Such extra features have not been detected in experiment, likely due to application of higher bias currents. Future experiments could be dedicated to the search of additional frustration features.. Overall, the ’frustrated frustration’ pattern ($f+f^{\prime}\in\mathbb{Z}$) is more stable than regular frustration. For an intuitive understanding, consider the equilibrium ($I=0$) configurations of loop currents  [^4] Our definition of loop currents, as shown in Fig. 3 h), is related to, but distinct from, the more commonly considered vortex picture. flowing through single plaquettes, see Fig. 3 h). Thus we find that, e.g., the resistance dip at $f\approx 2/3,f^{\prime}\approx 1/3$ resembles much more closely the configuration of $f=f^{\prime}=1/2$ than the one at $f=f^{\prime}=1/3$ – however, with an overall reduced loop current magnitude, indicating increased stability. Stability considerations of the various frustration features can also be understood in terms of the regular vortex model [55].
Conclusion. We demonstrated an array made with 4TJJs. For this we introduced an in-situ fabrication technique for arrays. A frustration pattern was measured which differs from the expectation for ordinary 2TJJ arrays. The difference could be explained by theoretically introducing a checker board lattice with two alternating flux patterns $f$ and $f^{\prime}$. The periodicity of the beating pattern can be connected to the area ratios $A$/$A^{\prime}$. In particular, while dips in the dc resistance can be linked to the sum $f+f^{\prime}$ being integer, the in general irrational area ratio leads to an irrational repartition of this total flux into the alternating plaquettes. This allows for the stabilization of the superconducting BKT phase at irrational fractions of flux penetrating individual plaquettes. This feature allows us to estimate the spatial expansion of the 4TJJ central weak link region, the result of which is compatible with the junction geometry. Overall, the here considered setup opens up a previously unexplored class of array system with alternating flux textures. Given that detailed understanding of vortex lattice structures and their dynamics are still an open question for regular lattices with a single flux $f$ [2], the here introduced alternating $f,f^{\prime}$ lattice will likely open up a new research direction in its own right (e.g., exploring the impact of capacitive quantum fluctuations on the phase diagram as in Ref. [8], now including incommensurate vortex features).
Acknowledgements. We thank Herbert Kertz for technical assistance. All samples have been prepared at the Helmholtz Nano Facility [60]. This work is funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany’s Excellence Strategy – Cluster of Excellence Matter and Light for Quantum Computing (ML4Q) EXC 2004/1 – 390534769, by the German Federal Ministry of Education and Research (BMBF) via the Quantum Futur project ‘MajoranaChips’ (Grant No. 13N15264) and the Quantum Futur project with grant no. 13N14891 (RR), as well as the Bavarian Ministry of Economic Affairs, Regional Development and Energy within Bavaria’s High-Tech Agenda Project ”Bausteine für das Quantencomputing auf Basis topologischer Materialien mit experimentellen und theoretischen Ansätzen” (grant no. 07 02/686 58/1/21 1/22 2/23).
SUPPLEMENTAL NOTE 1: Fabrication procedure
The stencil lithography fabrication process of the four-terminal Josephson junction (4TJJ) array with molecular beam epitaxy is described in the following. As a first step, a Si(111) substrate is covered by a 300-nm-thick SiO2 layer followed by a 100-nm-thick Si3N4 layer. Subsequently, the stencil mask is patterned into the Si3N4 layer by electron beam lithography and reactive ion etching. As a next step, the stencil mask is under-etched by hydrofluoric acid, removing the SiO2 layer below and revealing the Si(111) substrate surface. The stencil mask can be seen in main text Fig. 1 c). After inserting the pre-patterned substrate into an ultra high vacuum growth chamber, a 10-nm-thick Al2O3 interlayer is first deposited under rotation. The rotation ensures that the junction array area is completely covered by Al2O3. Next, a 20-nm-thick Pt layer is deposited under rotation. Here, the deposition at different angles due to the substrate rotation results in the deposition of Pt connecting the openings of the stencil mask. Finally, a 50-nm-thick superconducting Nb layer is deposited without rotation, resulting in a Nb electrode pattern defined by the opening of the stencil mask. As a last step, a 5-nm-thick Al2O3 capping layer is deposited under rotation. In main text Fig. 1 c), a scanning electron microscope image of the final device structure is shown. The four-terminal Josephson junction consists of four superconducting Nb islands connected at their tips by a central Pt weak link area. After the material deposition, the stencil mask is removed by applying ultrasonic power. The square unit cell of the array formed by a 4TJJ in each corner is indicated in main text Fig. 1 c) as well. To prevent screening of magnetic fluxes by screening currents in superconducting material surrounding the array, the Nb layer outside the device is removed by reactive ion etching (see suppl. section 7).
SUPPLEMENTAL NOTE 2: Measurement setup
The measurements have been performed in a dilution refrigerator with a base temperature of 70 mK equipped with a superconducting magnet. For measurements with magnetic fields below 80 mT, an external high-resolution current source was used to supply the bias current of the magnet. The voltage signal, from which the resistance ($R_{\text{dc}}$) and differential resistance ($dV/dI$) has been calculated, was measured with a quasi 4-point setup by applying a dc bias to the array. In additional measurements (suppl. section 9), the dc signal was superimposed with a smaller AC signal in order to measure the differential resistance using the lock-in technique and the ac signal has been used alone for ac resistance measurements dependent on magnetic field.
SUPPLEMENTAL NOTE 3: Second 30$\times$30 4TJJ array
The critical current oscillates with magnetic field in the same way as for the main text array. With a DC bias of 30 $\mu$A, that behavior leads to an oscillation of the resistance with magnetic field [see Fig. S4 a)]. Similar to the main text device, the frustrated frustration pattern is present. The damping of the oscillations of resistance and critical current can be reproduced.
At zero magnetic field and a temperature of 80 mK, the critical current is comparable to the main text device with $I_{\mathrm{c}}=59.75\;\mu$A [see Fig. S1 c) and d)]. The differential resistance is lower with $R=1.75\;\Omega$ close to the maximum applied current of 100 $\mu$A.
Supplemental Figure S1: Data of the second 30$\times$30 four-terminal Josephson junction array. a) Differential resistance-bias current-magnetic field diagram. The oscillations of the critical current are clearly visible. b) Scanning electron microscopy image of the device. c) I-V measurement at 80 mK. d) Differential resistance calculated from the data in c).
SUPPLEMENTAL NOTE 4: Reference two-terminal Josephson junction
On the chip with the presented arrays, a reference two-terminal Josephson junction has been fabricated in the same way as the arrays (see Fig. S2). The weak link length was determined to be 92 nm and its width is 920 nm. Although only a width of 809 nm is indicated in Fig. S2, Pt and Nb material under the edge of the bottom Si3N4 layer needs to be taken into account. The angle between the substrate surface and Nb beam was 57.7. The 300 nm-thick-SiO2 layer right at the edge of the trench is etched away, just like under the mask itself. This leaves the 100 nm-thick-Si3N4 layer under-etched at the edge and gives additional substrate space for the niobium shadow to be deposited. The extent of the Nb shadow under the Si3N4 edge can be derived from the top shadow of Fig. S2 (174 nm). The bottom shadow has a smaller distance from its edge than the top shadow because it is delimited by the lower edge of the Si3N4 layer (instead of the upper edge). That means that the bottom shadow will be $100\;\text{nm}/\text{tan}(57.7^{\circ})=63\;\text{nm}$ shorter than the top shadow. This results in a total bottom shadow of 111 nm, which has already been added to the width specification of this Josephson junction above (920 nm).
In Fig. S2, the junction is shifted horizontally with respect to the stencil mask [by $174\;\text{nm}\cdot\text{tan}(42.2^{\circ})=158\;\text{nm}$] because latter was rotated against the Nb beam during deposition.
Figure S3 shows the voltage and differential resistance dependent on the current. The device has a critical current of $I_{\text{c}}=56.51\;\mu$A and a differential resistance of $R=1.21\;\Omega$. The Josephson energy of the junction is $E_{\mathrm{J}}=\Phi_{0}I_{\mathrm{c}}/2\pi=116$ meV. Using a relative permittivity of one (because platinum is a metal) and the geometrical values given above (length of a SC contact is 600 $\mu$m), the coplanar junction capacitance is $C=5.6\cdot 10^{-17}\;$.
[…]
The Stewart-McCumber parameter for these values (assuming $R_{N}\approx R$) is $\beta_{\text{c}}=2eI_{\text{c}}R_{\text{N}}^{2}C/\hbar\approx 1.4\cdot 10^{-5}$. Therefore, the junction is in the over-damped regime. Although the junction is over-damped, the current-voltage characteristic of the device shows hysteretic behavior [see Fig. S3 a)]. This can be explained by an increase in electron temperature in the resistive state [1].
[…]
SUPPLEMENTAL NOTE 5: Magnetic hysteresis of the resistance pattern
@@
In the main text, we remove the hysteresis by showing only one sweep direction of the magnetic field, namely from negative to positive, and shift the data by 0.9 mT to the negative side.
The period of the Fraunhofer effect for the reference two-terminal Josephson junction (2TJJ) in Fig. S4 c) does not fit to the weak link area given in suppl. section 4. Possible reasons for this are an unequal current distribution, flux focusing, or the presence of Abrikosov vortices.
@@
Supplemental Figure S4: Resistance dependent on magnetic field for: (a) the main text 4TJJ array, (b) the second 4TJJ array, (c) and the reference two-terminal Josephson junction. The blue lines indicate a magnetic sweep direction from negative to positive and the orange lines from positive to negative. All devices show hysteretic behavior with magnetic field. The two upper resistance oscillations stem from the frustration pattern of the array, the lower resistance oscillation stems from the Fraunhofer effect of the two-terminal Josephson junction.
SUPPLEMENTAL NOTE 6: Determination of the resistance oscillation period
For the determination of the oscillation frequencies, fast Fourier transforms have been applied to the data in Fig. 2 b) and c) of the main text (see Fig. S5). The frequency spectra of both data curves match. The dominant frequency (period) is 0.16 mT-1 (6.25 mT). The period of the beating pattern cannot be seen in the spectrum, probably because it is too close to the zero peak.
Supplemental Figure S5: Fast Fourier transforms of the data curves in main text Fig. 2 b) (fine sweep) and c) (wide range). The dominant magnetic field frequency (period) is 0.16 mT-1 (6.25 mT). The other peaks of the spectrum are higher harmonics of that frequency.
SUPPLEMENTAL NOTE 7: Description of the array device
This section gives a detailed description of the array device in the main text, including geometrical properties of the array lattice (see Fig. S6). According to Fig. S6 a), the unit cell of the array is a square of size $A_{\text{UC}}=573\;\text{nm}\cdot 567\;\text{nm}=(570\;\text{nm})^{2}$. The geometry of the weak link area between two neighboring leads of one 4TJJ [see Fig. S6 b)] is described, on average, by the following dimensions: The width of the weak link is 129 nm and its length is 97 nm. The superconducting arms in the 4TJJ array have a length of 471 nm.
During fabrication of the device, a trench is etched into the SiO2 (300 nm) / Si2N3 (100 nm) layer, down to the Si(111) substrate. On the substrate, the device is fabricated and, afterwards, the whole chip is covered with superconducting material. To minimize disturbance effects, the superconductor around the device is etched away as a last fabrication step, leaving an edge of superconductor of approximately 1 µm around the trench behind. To prevent this closed loop of superconductor from generating screening currents, the edge is disconnected close to the ends of the device contacts [see Figs. S6 c) and d)].
@@
SUPPLEMENTAL NOTE 8: Characteristic resistively capacitively shunted junction (RCSJ) values between two leads of a 4TJJ in the arrray
In this section, relevant parameters for the theoretical description of the array are derived from geometrical considerations of the prior section, as well as from measurement data of the reference two-terminal Josephson junction. We assume that critical current density and resistivity, derived in suppl. section 4, are geometry independent. We then use these values, together with the 4TJJ geometry, to derive capacitance, resistance, and critical current between two leads of one 4TJJ.
The coplanar capacitance is calculated using the geometrical values given in suppl. section 7. The relative permeability $\epsilon_{r}$ is assumed to be one, since the weak link is a metal. This results in a capacitance of $C=2.729\cdot 10^{-18}$ F. The critical current between two 4TJJ leads is $I_{\mathrm{C}}=7.9$ µA. This results in a Josephson energy of $E_{\mathrm{J}}=\Phi_{0}I_{\mathrm{c}}/2\pi=16.2$ meV. The resistance is equal to $R=9.1\;\Omega$. The Stewart-McCumber parameter for these values (assuming $R_{N}\approx R$) is $\beta_{\text{c}}=2eI_{\text{c}}R_{\text{N}}^{2}C/\hbar\approx 5.4\cdot 10^{-6}$. Therefore, the junction is in the over-damped regime.
Although the junction is over-damped, the current-voltage characteristic of the array shows hysteretic behavior [see Fig. S7 a)]. This can be explained by an increase in electron temperature in the resistive state [1].
SUPPLEMENTAL NOTE 9: Additional measurements of the main text 30$\times$30 4TJJ array
@@
Supplemental Figure S7: IV measurements of the array discussed in the main text at 80 mK. a) The DC IV curve with a critical current of $I_{\mathrm{c}}=57\;\mu$A. b) AC differential resistance measured with the lock-in technique. The AC bias was 200 nA.
@@
Figure S9 shows the same resistance oscillations with magnetic field as before but, here, they were measured with the AC lock-in technique and no DC bias. The resulting resistance is lower than with the DC measurement because the time-dependent sinusoidal signal includes the superconducting state of the device as well. When comparing Fig. S9 a) with Fig. S4 a), one can see that the oscillations are identical in magnetic field, although the peak pattern is more symmetrical with the AC method.
@@
Figure S10 a) shows resistance measurements dependent on magnetic field and DC current. At higher currents, superconductivity is not realized at any magnetic field and the oscillation amplitudes decline.
Figure S10 b) shows resistance measurements dependent on magnetic field and temperature. At higher temperatures, superconductivity is not realized at any magnetic field and the oscillation amplitudes decline.
@@
SUPPLEMENTAL NOTE 10: Details of the resistively capacitively shunted junction model calculations
@@
We here summarize the model calculations for the theoretical discussion of the main text. We describe the multiterminal Josephson junction array in the overdamped regime (discarding capacitive contributions), we get at each node $j=(j_{x},j_{y})$ (where $j_{x,y}=1,\ldots,J_{x,y}$) the equation of motion,
$\frac{1}{2eR}\sum_{\langle j^{\prime}\rangle_{j}}\left(\dot{\phi}_{j}-\dot{\phi}_{j^{\prime}}\right)=-I_{c}\sum_{\langle j^{\prime}\rangle_{j}}\sin(\phi_{j}-\phi_{j^{\prime}}+a_{j,j^{\prime}})+\sum_{\langle j^{\prime}\rangle_{j}}\eta_{j,j^{\prime}}(t)\ ,$ (S1)
with parallel shunt resistance $R$ and critical current $I_{c}$. The vector potential $a_{j,j^{\prime}}$ accounts for the magnetic flux piercing the plaquettes of the array (which will be defined in detail futher below). While we assume zero temperature in the device for the discussion in the main text, we here include for completeness the current noise term $\eta_{j,j^{\prime}}$ due to thermal fluctuations from the resistor connecting nodes $j$ and $j^{\prime}$, where
$\langle\eta_{i,i^{\prime}}(t)\eta_{j,j^{\prime}}(t^{\prime})\rangle=\frac{2k_{B}T}{R}\delta_{i,j}\delta_{i^{\prime},j^{\prime}}\delta(t-t^{\prime})\ .$ (S2)